OpenStudy (anonymous):

help me !!! Nakim simplified 3 times the square root of 2x plus x times the square root of 8x minus 5 times the square root of 18x and got -10x times the square root of 2x for an answer. Part 1: Using complete sentences, explain what Nakim did wrong. Part 2: Show all your work to simplify the expression. (You can use "sqrt()" to show a square root. For example, 7 times the square root of a can be written as 7sqrt(a). )

OpenStudy (anonymous):

\[\sqrt{a}*\sqrt{b} = \sqrt{ab}\] \[3\sqrt{2x} + x*\sqrt{8x} - 5\sqrt{18x} = 3\sqrt{2x} + x*\sqrt{4}*\sqrt{2x} - 5\sqrt{9}*\sqrt{2x}\]

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

\[= 3\sqrt{2x} +2x*\sqrt{2x} -15\sqrt{2x} = (2x - 12)\sqrt{2x}\] not sure what Nakim did RIGHT lol. :P

OpenStudy (anonymous):

ok thank you so much

OpenStudy (anonymous):

glad i could help :)